# Particles before symmetry

### 05 Mar 2026 v5

@@ Abstract
The Standard Model of particle physics is usually cast in symmetry-first terms. On this approach, one begins with a symmetry group and postulates matter fields as objects transforming under its representations, without requiring that the group be grounded in, or derived from, independent geometric structures. Recently, a geometry-first formulation has been proposed, in which the relevant symmetries are not fundamental. In this paper I extend this approach to two central mechanisms of the Standard Model: spontaneous symmetry breaking and the Yukawa coupling, both essential for particles to acquire mass. These reformulations offer alternative explanations cast in purely geometric terms. In particular, the quantisation of charge arises here as a purely geometric consequence of the tensorial construction of matter fields from the fundamental bundles—a mechanism that is both more general and more transparent than the usual topological account based on the compactness of symmetry groups. More generally, I argue that a symmetry-first account in terms of principal and associated bundles admits a genuine geometry-first counterpart only under certain strict conditions.
@@ 1 Introduction
Mathematically equivalent formulations of a theory can differ profoundly in what they suggest. Feynman made the point in his Nobel Lecture:
The subsequent career of Feynman’s path integrals—equivalent to Schwinger’s operator methods, yet far more fertile—bears this out. So does Minkowski’s 1908 recasting of special relativity: it changed no predictions, but supplied the geometric ontology without which general relativity could hardly have been conceived.Theory-reformulation is a live philosophical topic, with positions ranging from deflationary instrumentalism to inflationary fundamentalism (\@BBOPcitep\@BAP\@BBN(Hunt, 2025)\@BBCP). [^2] According to Hunt (p. 5, ibid), instrumentalism treats reformulation as “merely a different choice of convention, no different in kind than a change in notation”; fundamentalism “proposes a metaphysical picture similar to David Lewis’s posits that some properties belong to an elite set of perfectly natural properties, with physics aiming to provide a partial inventory of these”. What I offer here should interest anyone who is not a thoroughgoing instrumentalist.
The reformulation I propose applies to classical gauge theory as it bears on the Standard Model—less radical than Feynman’s or Minkowski’s, but nonetheless fruitful, as I will argue. [^3] It does not apply to all gauge theories: this will matter in Section 5. The reformulation does not begin with symmetry so I call it geometry-first.
A knowledgeable reader will protest that gauge theory is already geometrical: principal fibre bundles and connections are, after all, the geometer’s stock-in-trade. True—but the usual formulation also builds symmetry into its foundations, and with it, an ontology that extends beyond the spaces where matter fields actually live.
Here is what I mean. In particle physics, matter fields are sections of vector bundles. The symmetry-first formulation adds principal fibre bundles on top, positing a structure group that coordinates the various matter sectors. I propose to dispense with this extra layer. If principal bundles appear at all in a geometry-first formulation, they must supervene entirely on the vector bundles—not the other way around.
@@
But dispensed with it can be. A geometry-first formulation has recently been proposed in which symmetries are not postulated and principal fibre bundles are unnecessary (\@BBOPcitep\@BAP\@BBN(Gomes, 2024, 2025b)\@BBCP). Symmetry groups become implicit: they arise as automorphism groups of fundamental vector bundles. The formulation is available only for gauge groups that are linear, and for representations built from the fundamental representation via tensor products, symmetrisation, and the like. It therefore demands an alignment between symmetry and geometry that the symmetry-first picture does not guarantee.
Even when the two formulations are mathematically equivalent, they come with different ontologies. For the Standard Model, the geometry-first picture posits three fundamental vector bundles over spacetime. Matter fields are sections of tensor products of these bundles. There is no separate space encoding the principal connection.
Change the formulation, and the explanations change with it. Four examples, developed below, show how.
First, in a non-Abelian vacuum Yang–Mills theory with Lie group $G$, the fundamental dynamical object is no longer a connection $\varpi$ on a $G$-principal bundle (or its spacetime representative $A_{\mu}^{I}$), but the covariant derivative $\nabla_{\mu}$ on a vector bundle whose automorphism group is $G$. This remains true even when no matter fields are present: the affine structure can be dynamical all by itself. Quantum numbers become geometric labels—the internal space a particle inhabits, and its tensor type within that space. (These ideas were explained in \@BBOPcitep\@BAP\@BBN(Gomes, 2024)\@BBCP; this paper focuses on the next three.)
Second, once symmetry groups drop out of the ground level, ‘symmetry-breaking’ requires reinterpretation. Vector bosons are replaced by covariant derivatives of fundamental bundles, which are not on the same footing as matter fields. How could they ‘acquire mass’ in the usual sense?
Third, Yukawa couplings. In the symmetry-first formulation, Yukawa terms are scalars formed from sections of different associated bundles, requiring explicit bridges between representation spaces—bridges that need not be unique. In the geometric picture, these scalars arise through inner products and contractions that are already determined by the bundle structures.
Fourth, the quantisation of charge. The symmetry-first picture attributes it to the topology of the compact group $U(1)$. The geometric picture derives it from the discrete algebraic structure of tensor powers of the fundamental bundles. This means that even for non-compact groups, where symmetry would permit continuity, geometry enforces discreteness.
Under \@BBOPcitet\@BAP\@BBNHunt (2025)\@BBCP’s lights, such a reformulation is non-trivial: it suggests a different ontology, offers new explanations, and displays an epistemic difference in how it solves problems.444“Successful reformulations clarify what we need to know to solve problems, improving our understanding of the world” (ibid. p. 5); “Within [the] shared domain of problems [that they both solve], significant reformulations display an epistemic difference, while trivial reformulations do not” (ibid. p. 9).
@@
A word of caution about scope. The VB-POV applies to a proper subclass of gauge theories: those whose gauge groups arise as automorphism groups of structured vector spaces, with matter built tensorially from fundamental bundles. As a theoretical framework, this is genuinely narrower than the PFB-POV—it excludes models that the principal-bundle formulation can perfectly well accommodate. But the Standard Model lies within the overlap, and within that overlap the two formulations are empirically equivalent. What changes is which features of the physics appear primitive and which derived. Two kinds of claim should therefore be kept apart throughout: claims about explanatory structure—the Higgs and Yukawa reformulations of Sections 34—and claims about admissible theory-space—notably the slack diagnosis and the charge-quantisation discussion of Section 5. The former are reformulations in the strict sense; the latter amount to extra constraints that I argue physicists already tacitly impose.
@@ 2.1 Gauge theory and principal fibre bundles: the symmetry-first formulation
Figure 1: The principal $G$-bundle , with structure group $G$, over the manifold $M$, with a principal connection $\varpi$ (a $\mathfrak{g}$-valued one-form on $P$), abbreviated by $P(M,G,\omega)$, and its associated vector bundles $E_{i}:=P\times_{\rho_{i}}V_{i}$, where $\rho_{i}:G\rightarrow V_{i}$ is a representation of the Lie group—determined by a particle’s quantum numbers—onto the vector space representing the typical fibre, $V_{i}$ which is linearly isomorphic to $\pi_{i}^{-1}(x)$, for $x\in M$ and $\pi_{i}:E\rightarrow M$ the projection of the vector bundle onto its base space (spacetime). The covariant derivatives $\nabla_{i}$ are the ones induced by $\varpi$, as per Equation (A.6). See Appendix A for more details.
@@
Where the last inequivalence holds iff $\rho_{1}(g)\rho_{2}^{-1}(g)\neq\mathbb{1},\,\,\forall g$, i.e. the equivalence holds iff $\rho_{1}\neq\rho_{2}$. Thus we find that for the map (2.4) to be well-defined, we must have $\rho_{1}=\rho_{2}$.
Indeed, in physics, we are often faced with situations in which $E_{1}$ and $E_{2}$ have the same typical fibre, are associated to the same group, and yet have different representations. A simple example is when one of the representations is the trivial, or singleton, one, and the other is the fundamental (or any other). [^7] A slightly more sophisticated example is as follows. Let $G=U(1),V=\mathbb{C}^{k}$, and $\rho_{i}=n_{i}$, which acts as $e^{in_{i}}\mathbb{1}$ on $\mathbb{C}^{k}$. Then for $n_{i}\neq n_{j}$ for $i\neq j$ the map (2.4) is not well-defined, as can easily be verified. This occurs many times in the Standard Model: for fermions to acquire mass, one must relate sections of bundles that have different representations, since they represent different particles.
@@ 2.2 Gauge theory and vector bundles: the geometry-first formulation
Matters of gauge invariance are now also seen under a different lens. Any composite object constructed from the basic dynamical variables—the $\nabla_{n}$ and other tensor fields—will be tensorial, that is, covariant under the corresponding automorphism groups; and any scalar formed from such quantities will be invariant under those groups. In this respect, the description closely parallels that of classical general relativity in modern treatments employing abstract-index notation: such treatments scarcely mention “gauge invariance” or “coordinate invariance”; all they require is that covariance be secured at the ground level. [^11] Upon quantisation, as I argue in \@BBOPcitep\@BAP\@BBN(Gomes, 2025a)\@BBCP, superpositions of states may require relating objects across distinct classical possibilities. That, I contend, is where “gauge fixing”—or, more broadly, what I call representational schemes—enters. Gauge fixings become necessary when a fixed reference across classical states is needed. If such references are physical, they can, incidentally, be used to describe content in a gauge-invariant (or gauge-fixed) manner, in the traditional PFB-POV sense.
@@
$\iota_{\xi}\Omega(\xi_{1},\cdots,\xi_{m-1})=\Omega(\xi,\xi_{1},\cdots,\xi_{m}),$ (2.9)
@@
One might object that a parallel representation-theoretic argument could be mounted. Perhaps there are general ways to relate arbitrary representation spaces that mirror the geometric ones. That may be true, but it is beside the point. The virtue of the geometric route is that it trades purely on geometrical language, speaking directly to those trained in geometry rather than representation theory. The mere availability of a formulation that sidesteps algebraic machinery is already a win.
@@
$\big(E^{3},M,\mathbb{C}^{3},\langle\cdot,\cdot\rangle_{3},\epsilon_{3}\big),\qquad\big(E^{2},\mathbb{C}^{2},\langle\cdot,\cdot\rangle_{2},\epsilon_{2}\big),\qquad\big(E^{1},\mathbb{C},\langle\cdot,\cdot\rangle_{1}\big).$ (2.10)
@@ 3 The Higgs mechanism in the geometry-first formulation
In the standard presentation (see, e.g., \@BBOPcite\@BAP\@BBN(Tong, 2025, Ch. 2)\@BBCP\@BBOPcite\@BAP\@BBN(Bleecker, 1981, Ch. 10.3)\@BBCP), the Higgs mechanism is usually described in the language of spontaneous symmetry breaking. One is then compelled to invoke stabilisers, quotients of groups, Killing forms on Lie algebras, Goldstone’s theorem, among other tools. In this section I outline an equivalent treatment, differing only in that it is phrased entirely in the language of vector bundles; the essential structure becomes visible without appeal to symmetry. It serves as a proof of principle that, in most applications, rerouting our demonstrations to avoid talk of symmetry does not consign us to intractably long alternatives (a common objection to some well-known reformulations; e.g. the Einstein algebra formulation of general relativity \@BBOPcitep\@BAP\@BBN(Geroch, 1972)\@BBCP).
@@ 3.1 The Higgs field as a background
where the caligraphic $\mathcal{L}$ will denote Lagrangian densities.
@@
First, note that $\nabla e_{0}$ does not contain all the information in $\nabla$: the components of $\nabla$ that do not enter $\nabla e_{0}$ are destined to remain ‘massless’. From (3.4), only some of the components of $\nabla$ that rotate $e_{0}$ into directions orthogonal to $e_{0}$ appear in the quadratic term, while those preserving $e_{0}$ drop out. In more detail, from (3.4), using $\mathsf{Re}\,\langle\,.\,$$\rangle$ (see footnote 14), we project $\nabla e_{0}$ along $e_{0}$ and its orthogonal complement to find a geometric description of (3.5) as follows.
Viewing the unit section $e_{0}$ as defining a rank-one subbundle $L:=\mathrm{span}\{e_{0}\}\subset E$, we can interpret its orthogonal complement with respect to $\text{Re}\langle\cdot,\cdot\rangle$ $L^{\perp}$ as a normal bundle to $L$ within $E$. The covariant derivative $\nabla e_{0}$ then defines a one-form with values in $L^{\perp}$; it probes how $L^{\perp}$ curves within $L$. Namely, what we can call a mixed shape operator for $L^{\perp}$: [^15] In differential geometry, given a hypersurface $S\subset M$ with unit normal $n$, the standard shape operator for $S$ is defined as a tensorial operator $K_{n}:TS\rightarrow TS$, namely $K=\nabla n$. Here our operator will mix indices, $K_{n}:TM\rightarrow L^{\perp}$, but it can still be thought of as an operator describing the shape of $L^{\perp}$ within $L$ as one goes around $M$.
@@
This concludes the classical, qualitative, account of ‘mass acquisition’ for the affine structure using the Higgs as a fixed background structure. Two points of contact with the standard formalism should be noted. First, ‘mass acquisition’ is, at this stage, a statement about the quadratic form (3.8), not yet about gauge bosons in the particle-physics sense. The identification of specific massive and massless components requires writing $\nabla={\mathrm{d}}+\omega$ in a frame adapted to $e_{0}$; this is carried out in Appendix B. Second, the decomposition above is local: it presupposes that $\varphi$ is nowhere vanishing. On nontrivial backgrounds the global picture requires patching, but the geometric content of (3.8) is unaffected.

### 23 Mar 2026 v6

@@ Abstract
The Standard Model of particle physics is standardly formulated in terms of principal fibre bundles and their associated representations—what may be called a symmetry-first approach. This paper develops an alternative geometry-first formulation in which the fundamental objects are Hermitian vector bundles equipped with compatible covariant derivatives, and symmetry groups arise as automorphism groups rather than being postulated. Working entirely in this framework, I give new derivations of two central mechanisms of the Standard Model. First, the Higgs mechanism: mass acquisition for gauge bosons is identified with the extrinsic curvature of a sub-bundle singled out by the Higgs vacuum, requiring no appeal to symmetry breaking, Goldstone’s theorem, or gauge fixing. Second, the Yukawa mechanism: fermion mass terms are constructed as fibrewise contractions determined by the inner-product and orientation structures of the fundamental bundles, eliminating a centraliser ambiguity that arises in the standard representation-theoretic treatment. I also show that charge quantisation follows from the discrete algebraic structure of tensor powers of the fundamental line bundle—a derivation that applies even when the relevant automorphism group is non-compact. Finally, I identify the structural conditions under which the symmetry-first and geometry-first formulations are genuinely equivalent, showing that the Standard Model satisfies these conditions but that theories based on exceptional Lie groups do not.
@@ 1 Introduction
This paper develops a reformulation of classical gauge theory, as it bears on the Standard Model, in which the fundamental mathematical objects are vector bundles rather than principal bundles. In the standard formulation—which I call symmetry-first—one begins with a structure group $G$ acting on a principal bundle and constructs matter fields as sections of associated vector bundles. In the reformulation developed here—which I call geometry-first—one begins instead with Hermitian vector bundles equipped with compatible covariant derivatives, and the group $G$ appears only as the automorphism group of the fibre. [^1] The geometry-first formulation does not apply to all gauge theories: this will matter in Section 5. For the Standard Model gauge group $SU(3)\times SU(2)\times U(1)$, the two formulations yield the same physics, but they are not mathematically equivalent: they differ in mathematical starting point, in the explanations they offer, in the data they record, and in the class of theories they naturally accommodate (see \@BBOPcitep\@BAP\@BBNGomes, 2026\@BBCP for a categorical analysis).
A knowledgeable reader will protest that gauge theory is already geometrical: principal fibre bundles and connections are, after all, the geometer’s stock-in-trade. True—but the usual formulation also builds symmetry into its foundations, and with it a layer of structure that extends beyond the spaces where matter fields actually live. In particle physics, matter fields are sections of vector bundles. The symmetry-first formulation adds principal fibre bundles on top, positing a structure group that coordinates the various matter sectors. I propose to dispense with this extra layer. If principal bundles appear at all in a geometry-first formulation, they must supervene entirely on the vector bundles—not the other way around.
@@
But dispensed with it can be. A geometry-first formulation has recently been proposed in which symmetries are not postulated and principal fibre bundles are unnecessary (\@BBOPcitep\@BAP\@BBN(Gomes, 2024, 2025a)\@BBCP). Symmetry groups become implicit: they arise as automorphism groups of fundamental vector bundles. The formulation is available only for gauge groups that are linear, and for representations built from the fundamental representation via tensor products, symmetrisation, and the like. It therefore demands an alignment between symmetry and geometry that the symmetry-first picture does not guarantee. For the Standard Model, the geometry-first picture posits three fundamental vector bundles over spacetime. Matter fields are sections of tensor products of these bundles. There is no separate space encoding the principal connection.
The present paper extends this programme to two mechanisms central to the Standard Model: the Higgs mechanism and the Yukawa coupling, both essential for particles to acquire mass. It makes four main contributions.
Higgs mechanism. I derive the classical Higgs mechanism entirely in vector-bundle language (Section 3). A non-vanishing Higgs section singles out a rank-one sub-bundle $L\subset E$; the covariant derivative of the unit Higgs section defines a shape operator (extrinsic curvature) for the orthogonal complement $L^{\perp}$. The quadratic dependence on the connection that characterises mass acquisition arises from this shape operator. Components of the connection preserving the Higgs direction remain massless—corresponding, in symmetry-first language, to the unbroken gauge group. No appeal to symmetry breaking, Goldstone’s theorem, or gauge fixing is needed. In the linearised regime, the Higgs itself acquires mass through the curvature of the Mexican-hat potential, again expressed without group-theoretic apparatus.
Yukawa mechanism. I reconstruct the Standard Model Yukawa couplings as fibrewise contractions on the fundamental bundles $E^{1},E^{2},E^{3}$ (Section 4). In the symmetry-first formulation, Yukawa terms are $G$-invariant maps between representation spaces, determined only up to the centraliser of $G$; the geometry-first formulation replaces these with inner products and contractions already present on the bundles. The quark Yukawa term is shown to depend essentially on the volume form $\epsilon$ of $\mathbb{C}^{2}$, providing a geometric reason for the appearance of $SU(2)$ rather than $U(2)$.
Charge quantisation. I show that the quantisation of electric charge follows from the discrete algebraic structure of tensor powers of the fundamental line bundle $E^{1}$ (Section 5). This derivation does not rely on the compactness of $U(1)$: even if the automorphism group were non-compact, the tensorial construction would enforce discrete charges.
Structural comparison. I identify the conditions under which a symmetry-first formulation in terms of principal and associated bundles admits a genuine geometry-first counterpart (Section 5). The geometry-first picture requires each factor of $G$ to arise as the automorphism group of a fundamental bundle—a condition that structures the relationship between group and geometry more tightly than the principal-bundle picture demands. The Standard Model satisfies these conditions; theories based on exceptional Lie groups such as $E_{8}$ do not. Even on the overlap, the two formulations are not mathematically equivalent: the principal-bundle formulation does not record the tensorial genealogy of matter bundles from fundamental ones, as shown in a companion paper \@BBOPcitep\@BAP\@BBN(Gomes, 2026)\@BBCP.
The geometry-first formulation changes no predictions of the Standard Model. Its interest lies in the alternative explanations it provides: in each case above, a result that the symmetry-first formulation explains by appeal to group representations receives an independent explanation in terms of the geometric structures of vector bundles. Whether one regards this as evidence for a different underlying picture or merely as a useful change of perspective is a question I will return to briefly in the conclusions; the bulk of the paper is concerned with the mathematical content itself.
@@
A word of caution about scope. The VB-POV applies to a proper subclass of gauge theories: those whose gauge groups arise as products of automorphism groups of structured vector spaces, with matter built tensorially from fundamental bundles (a fundamental bundle need not contribute to any matter bundle: vacuum Yang–Mills sectors are legitimate). As a theoretical framework, this is genuinely narrower than the PFB-POV—it excludes models that the principal-bundle formulation can perfectly well accommodate. But the Standard Model lies within the overlap, and within that overlap the two formulations are empirically equivalent. What changes is which features of the physics appear primitive and which derived—and, as shown in a companion paper \@BBOPcitep\@BAP\@BBN(Gomes, 2026)\@BBCP, the symmetry-first data do not determine the geometry-first starting point even when both are available. Two kinds of claim should therefore be kept apart throughout: claims about explanatory structure—the Higgs and Yukawa reformulations of Sections 34—and claims about admissible theory-space—notably the slack diagnosis and the charge-quantisation discussion of Section 5. The former are reformulations in the strict sense; the latter amount to extra constraints that I argue physicists already tacitly impose.
@@ 2.1 Gauge theory and principal fibre bundles: the symmetry-first formulation
Figure 1: The principal $G$-bundle , with structure group $G$, over the manifold $M$, with a principal connection $\varpi$ (a $\mathfrak{g}$-valued one-form on $P$), abbreviated by $P(M,G,\varpi)$, and its associated vector bundles $E_{i}:=P\times_{\rho_{i}}V_{i}$, where $\rho_{i}:G\rightarrow V_{i}$ is a representation of the Lie group—determined by a particle’s quantum numbers—onto the vector space representing the typical fibre, $V_{i}$ which is linearly isomorphic to $\pi_{i}^{-1}(x)$, for $x\in M$ and $\pi_{i}:E\rightarrow M$ the projection of the vector bundle onto its base space (spacetime). The covariant derivatives $\nabla_{i}$ are the ones induced by $\varpi$, as per Equation (A.6). See Appendix A for more details.
@@
Where the last inequivalence holds iff $\rho_{1}(g)\rho_{2}^{-1}(g)\neq\mathbb{1},\,\,\forall g$, i.e. the inequivalence holds iff $\rho_{1}\neq\rho_{2}$. Thus we find that for the map (2.4) to be well-defined, we must have $\rho_{1}=\rho_{2}$.
Indeed, in physics, we are often faced with situations in which $E_{1}$ and $E_{2}$ have the same typical fibre, are associated to the same group, and yet have different representations. A simple example is when one of the representations is the trivial, or singleton, one, and the other is the fundamental (or any other). [^4] A slightly more sophisticated example is as follows. Let $G=U(1),V=\mathbb{C}^{k}$, and $\rho_{i}=n_{i}$, which acts as $e^{in_{i}\theta}\mathbb{1}$ on $\mathbb{C}^{k}$. Then for $n_{i}\neq n_{j}$ for $i\neq j$ the map (2.4) is not well-defined, as can easily be verified. This occurs many times in the Standard Model: for fermions to acquire mass, one must relate sections of bundles that have different representations, since they represent different particles.
@@ 2.2 Gauge theory and vector bundles: the geometry-first formulation
Matters of gauge invariance are now also seen under a different lens. Any composite object constructed from the basic dynamical variables—the $\nabla_{n}$ and other tensor fields—will be tensorial, that is, covariant under the corresponding automorphism groups; and any scalar formed from such quantities will be invariant under those groups. In this respect, the description closely parallels that of classical general relativity in modern treatments employing abstract-index notation: such treatments scarcely mention “gauge invariance” or “coordinate invariance”; all they require is that covariance be secured at the ground level. [^8] Upon quantisation, as I argue in \@BBOPcitep\@BAP\@BBN(Gomes, 2025b)\@BBCP, superpositions of states may require relating objects across distinct classical possibilities. That, I contend, is where “gauge fixing”—or, more broadly, what I call representational schemes—enters. Gauge fixings become necessary when a fixed reference across classical states is needed. If such references are physical, they can, incidentally, be used to describe content in a gauge-invariant (or gauge-fixed) manner, in the traditional PFB-POV sense.
@@
$\iota_{\xi}\Omega(\xi_{1},\cdots,\xi_{m-1})=\Omega(\xi,\xi_{1},\cdots,\xi_{m-1}),$ (2.9)
@@
One might object that a parallel representation-theoretic argument could be mounted. Perhaps there are general ways to relate arbitrary representation spaces that mirror the geometric ones. That may be true, but it is beside the point. The virtue of the geometric route is that it trades purely on geometrical language, speaking directly to those trained in geometry rather than representation theory. The mere availability of a formulation that sidesteps algebraic machinery is already significant.
@@
$\big(E^{3},M,\mathbb{C}^{3},\langle\cdot,\cdot\rangle_{3},\epsilon_{3}\big),\qquad\big(E^{2},M,\mathbb{C}^{2},\langle\cdot,\cdot\rangle_{2},\epsilon_{2}\big),\qquad\big(E^{1},M,\mathbb{C},\langle\cdot,\cdot\rangle_{1}\big).$ (2.10)
@@ 3 The Higgs mechanism in the geometry-first formulation
In the standard presentation (see, e.g., \@BBOPcite\@BAP\@BBN(Tong, 2025, Ch. 2)\@BBCP\@BBOPcite\@BAP\@BBN(Bleecker, 1981, Ch. 10.3)\@BBCP), the Higgs mechanism is usually described in the language of spontaneous symmetry breaking. One is then compelled to invoke stabilisers, quotients of groups, Killing forms on Lie algebras, Goldstone’s theorem, among other tools. In this section I outline an equivalent treatment, differing only in that it is phrased entirely in the language of vector bundles; the essential structure becomes visible without appeal to symmetry. It also shows that avoiding explicit symmetry need not produce unwieldy proofs (a common objection to some reformulations; e.g. the Einstein algebra formulation of general relativity \@BBOPcitep\@BAP\@BBN(Geroch, 1972)\@BBCP).
@@ 3.1 The Higgs field as a background
where the calligraphic $\mathcal{L}$ will denote Lagrangian densities.
@@
First, note that $\nabla e_{0}$ does not contain all the information in $\nabla$: the components of $\nabla$ that do not enter $\nabla e_{0}$ are destined to remain ‘massless’. From (3.4), only some of the components of $\nabla$ that rotate $e_{0}$ into directions orthogonal to $e_{0}$ appear in the quadratic term, while those preserving $e_{0}$ drop out. In more detail, from (3.4), using $\mathsf{Re}\,\langle\,.\,$$\rangle$ (see footnote 11), we project $\nabla e_{0}$ along $e_{0}$ and its orthogonal complement to find a geometric description of (3.5) as follows.
Viewing the unit section $e_{0}$ as defining a rank-one subbundle $L:=\mathrm{span}\{e_{0}\}\subset E$, we can interpret its orthogonal complement with respect to $\text{Re}\langle\cdot,\cdot\rangle$ $L^{\perp}$ as a normal bundle to $L$ within $E$. The covariant derivative $\nabla e_{0}$ then defines a one-form with values in $L^{\perp}$; it probes how $L^{\perp}$ curves within $E$. Namely, what we can call a mixed shape operator for $L^{\perp}$: [^12] In differential geometry, given a hypersurface $S\subset M$ with unit normal $n$, the standard shape operator for $S$ is defined as a tensorial operator $K_{n}:TS\rightarrow TS$, namely $K=\nabla n$. Here our operator will mix indices, $K_{n}:TM\rightarrow L^{\perp}$, but it can still be thought of as an operator describing the shape of $L^{\perp}$ within $L$ as one goes around $M$.
@@
This concludes the classical, qualitative, account of ‘mass acquisition’ for the affine structure using the Higgs as a fixed background structure. Two points of contact with the standard formalism. First, ‘mass acquisition’ is, at this stage, a statement about the quadratic form (3.8), not yet about gauge bosons in the particle-physics sense. The identification of specific massive and massless components requires writing $\nabla={\mathrm{d}}+\omega$ in a frame adapted to $e_{0}$; this is carried out in Appendix B. Second, the decomposition above is local: it presupposes that $\varphi$ is nowhere vanishing. On nontrivial backgrounds the global picture requires patching, but the geometric content of (3.8) is unaffected.
The geometric picture also makes vivid a fact that the symmetry-first formulation tends to obscure. After symmetry breaking, the Higgs vacuum singles out a rank-one subbundle $L=\mathrm{span}\{e_{0}\}\subset E$, splitting $E\cong L\oplus L^{\perp}$. But there is no reason for the covariant derivative to preserve this splitting: $\nabla P_{L}\neq 0$ in general, where $P_{L}$ is the orthogonal projector onto $L$. In a local frame adapted to the splitting, the connection one-form need not be block-diagonal, and the off-diagonal components—precisely those that acquired mass—measure the failure of $L$ and $L^{\perp}$ to be parallel subbundles. Hence parallel transport around a closed loop can mix the two components. Applied to the electroweak theory (Section 4), this means that a left-handed electron, defined as the component of $\mathbf{L}_{L}$ along the Higgs direction, can return from a loop as a mixture of electron and neutrino. This is certainly known to experts, but it is not something one reads off from the unbroken $U(1)_{\mathrm{EM}}$ alone; in the symmetry-first presentation, the unbroken subgroup preserves the splitting by definition, and the mixing is hidden in the broken generators. The geometry-first picture makes it immediate: the splitting is a condition on the section, not on the connection. This remains true even on topologically non-trivial backgrounds: topology governs whether the splitting $E\cong L\oplus L^{\perp}$ exists globally, whereas its preservation is a further differential condition on the connection.